# Univariate-Guided Sparse Regression

### v4

@@ 1 Introduction
Here is an example, taken from unpublished work with collaborators of the third author. (Since the work is unpublished, we do not give background details). The dataset is from cancer proteomics, with 559 biomarkers measured on 81 patients, 20 healthy and 61 with cancer. We divided the data into 70% training and 30% test, and applied both the lasso and uniLasso to the training set. Both methods performed well, giving cross-validation and test set error of about 5%. Table 1 shows the results of lasso and uniLasso applied to this data, using cross-validation to choose the model. In the top table, 3 of the 9 markers change sign from the univariate to multivariate; marker 9 is especially worrisome, as it changes from 1.063 to -14.057! By design, uniLasso in the bottom table has no sign changes. Further, its univariate coefficients are much larger (in absolute value) than those of the lasso. This makes it a more credible model.
The remainder of this paper is organized as follows. Section 2 provides a detailed description of the UniLasso methodology. We examine the relationship between UniLasso and the Adaptive Lasso in Section 3. We compare UniLasso to least squares and Lasso in the well-known diabetes data set in Section 4: here we show the utility of unregularized uniLasso as an alternative to least squares. In Section 5 we look at the orthonormal design case, where we can derive explicit an expression for the UniLasso solution. Section 6 presents some theoretical results on support recovery and mean square error consistency. Section 7 presents simulation studies comparing UniLasso to Lasso and other benchmark methods. Section 8 we apply UniLasso to real-world datasets, and in Section 9 we present multiclass and GWAS examples. Section 10 explores the setting where additional data is available, not full data but just the univariate coefficients. In Section 11 we study whether cross-validation works properly in the UniLasso setting. We discuss UniLasso for GLMs and the Cox Survival Model, as well as computation in Sections 12 and 13 Lastly, Section 14 concludes with a discussion of future directions.
Coefficients in model chosen by Lasso
Biomarker Univariate LS coefficient Lasso coefficient
1 -2.061 -7.323
2 3.545 39.834
3 -2.152 -7.516
4 -0.182 -2.174
5 -6.284 -0.574
6 -1.717 -4.794
7 7.888 -6.103
8 1.063 -14.057
9 -2.435 6.156
Coefficients in model chosen by uniLasso
@@
A 6.819 0.501
B 14.854 26.756
C 6.757 1.962
D 11.170 13.827
E -6.284 -12.171
F 9.158 22.807
G 8.654 19.721
H 5.055 3.204
Table 1: Proteomics study. Top table: Univariate least squares coefficients and lasso coefficients, for model chosen by Lasso. Bottom table: same for model chosen by uniLasso. All biomarkers are different in the two tables, with the exception of #5 in the first table corresponding to E in the second.
@@ 2.2 UniLasso with no-regularization
[…]
We call this “UniReg’ for univariate-guided regression and it represents an interesting alternative to the usual least squares estimates. For example, the non-negative constraint still can provide sparsity even though there is no $\ell_{1}$ penalty. And it gives a unique solution even if $p>n$. The standard error and distribution of the estimated coefficients can be estimated via the bootstrap. We compare UniReg to least squares in the diabetes example of Section 4.
@@ 6 Theoretical analysis
Let $\delta_{j,k}$ denote the population value of the univariate regression of $X_{j}$ on $X_{k}$. Suppose that for every pair of distinct indices $j,k\in S$, $\delta_{j,k}\geq 0$ if $\gamma_{j}$ and $\gamma_{k}$ have the same sign, and $\delta_{j,k}\leq 0$ if $\gamma_{j}$ and $\gamma_{k}$ have opposite signs. Then $\beta_{j}$ is nonzero and has the same sign as $\gamma_{j}$ for each $j\in S$. Moreover, $|\theta_{j}|\geq|\gamma_{j}|$ for each $j\in S$.
Corollary
The statement of Theorem 2 is “if and only if”.
@@ 7 Some simulation results
8 Real data examples
@@ 9.1 Multiclass setting and small round blue cell cancers
mean(err/25) sd(err/25) mean(supp) sd(supp)
@@ 14 Discussion
A draft version of the UniLasso package can be found at
@@ Acknowledgments
The authors would like to thank Jerome Friedman, Alden Green, Sam Gross, Lucas Janson, Mert Pilanci, Jonathan Taylor, Ryan Tibshirani and Hui Zou for helpful discussions. We especially thank Ryan for posing and proving the Corollary to Theorem 2. S.C. was supported by NSF grants DMS-2113242 and DMS-2153654. T.H was supported by grant 5R01 EB 001988-21 from the National Institutes of Health; R.T. was supported by the NIH (5R01EB001988- 16) and the NSF (19DMS1208164).
@@
Proof of Corollary to Theorem 2.
According to Theorem 2,
$Y=\sum_{k\in S}\beta_{k}X_{k}+\varepsilon$ (28)
$\beta_{j}=\frac{\text{Cov}(Y,X_{j})}{\text{Var}(X_{j})}$ (29)
$\beta_{j}=\sum_{k\in S}Y_{k}\frac{\text{Cov}(X_{k},X_{j})}{\text{Var}(X_{j})}$ (30)
Defining
$s_{kj}=\frac{\text{Cov}(X_{k},X_{j})}{\text{Var}(X_{j})},$ (31)
we can write
$\beta_{j}=\sum_{k\in S}\frac{Y_{k}}{Y_{j}}s_{kj}.$ (32)
Let
$A_{j}=\left\{k\in S:\text{sign}\left(\frac{\delta_{k}}{\delta_{j}}\right)=\text{sign}(s_{kj})\right\}.$ (33)
When $s_{kj}=0$, we allow this sign to be arbitrary.
Thus, for $\frac{\beta_{j}}{\delta_{j}}>0$, it is necessary and sufficient that
$\sum_{k\in A_{j}}\frac{\delta_{k}}{\delta_{j}}s_{kj}+\sum_{k\notin A_{j}}\frac{\delta_{k}}{\delta_{j}}s_{kj}>0.$ (34)
Since the first sum corresponds to positive contributions and the second to negative contributions, we get:
$\sum_{k\notin A_{j}}\left|\delta_{k}\right||s_{kj}|\leq\sum_{k\in A_{j}}\left|\delta_{k}\right||s_{kj}|.$ (35)
Here, the left-hand side represents the total ”disagreement” contribution, while the right-hand side represents the total ”agreement” contribution

### v7

@@ 1 Introduction
Here is an example, taken from unpublished work with collaborators of the third author. (Since the work is unpublished, we do not give background details). The dataset is from cancer proteomics, with 559 biomarkers measured on 81 patients, 20 healthy and 61 with cancer. We divided the data into 70% training and 30% test, and applied both the lasso and uniLasso to the training set. Both methods performed well, giving cross-validation error of about 5%, and 4 and 3 errors respectively out of 30 test samples. Table 1 shows the results of lasso and uniLasso applied to this data, using cross-validation to choose the model. In the top table there are 2 sign changes from univariate to multivariate: #3 goes from mildly protective to a negative indicator, while # 7 goes in the opposite direction.
By design, uniLasso in the bottom table has no sign changes. Further, its univariate coefficients are much larger (in absolute value) than those of the lasso. This makes it a more credible and stable model. We investigate stability in Section 10.1.
The remainder of this paper is organized as follows. Section 2 provides a detailed description of the UniLasso methodology. We examine the relationship between UniLasso and the Adaptive Lasso in Section 3. We compare UniLasso to least squares and Lasso in the well-known diabetes data set in Section 4: here we show the utility of unregularized uniLasso as an alternative to least squares. In Section 5 we look at the orthonormal design case, where we can derive explicit an expression for the UniLasso solution. Section 6 presents some theoretical results on support recovery and mean square error consistency. Section 7 presents simulation studies comparing UniLasso to Lasso and other benchmark methods, while Section 8.1 examines uniReg, the unregularized ($\lambda=0$) case. Section 9 applies UniLasso to real-world datasets, and in Section 10 we present an “adversarial attack” view, and multiclass and GWAS examples. Section 11 explores the setting where additional data is available, not full data but just the univariate coefficients. In Section 12 we study whether cross-validation works properly in the UniLasso setting. We discuss UniLasso for GLMs and the Cox Survival Model, as well as computation in Sections 13 and 14. Lastly, Section 15 concludes with a discussion of future directions.
Model chosen by Lasso
Biomarker Univariate LS coefficient Lasso coefficient Sign change?
1 -12.709 -1.845
2 25.269 37.492
3 -0.34 6.245 X
4 -24.74 -11.262
5 -17.966 -11.908
6 -49.24 -18.919
7 1.81 -6.265 X
8 22.909 13.995
9 -37.646 -5.415
10 13.024 0.79
11 -10.141 -2.325
12 -22.383 -12.745
Model chosen by uniLasso
@@
Biomarker univariate multivariate
A 58.924 4.982
B 25.269 6.286
C -49.24 -13.667
D 73.844 38.448
E 69.152 23.034
F -54.94 -9.261
G 77.688 9.458
Table 1: Proteomics study. Top table: Univariate least squares coefficients and lasso coefficients, for model chosen by Lasso. Bottom table: same for model chosen by uniLasso. All biomarkers are different in the two tables, with the exception of #2 and #6 in the first table corresponding to B and C in the second.
@@ 2.2 UniLasso with no-regularization
${\rm minimize}_{\theta}\;\left\{\frac{1}{n}\sum_{i=1}^{n}\Bigl{(}y_{i}-\theta_{0}-\sum_{j=1}^{p}\theta_{j}{(\hat{\beta}_{0j}^{-i}+\hat{\beta}_{j}^{-i}x_{ij})}\Bigr{)}^{2}\right\}\quad\mbox{ with $\theta_{j}\geq 0\;\forall j$}.$ (4)
We call this “UniReg’ for univariate-guided regression and it represents an interesting alternative to the usual least squares estimates. For example, the non-negative constraint still can provide sparsity even though there is no $\ell_{1}$ penalty. And it gives a unique solution even if $p>n$. The standard error and distribution of the estimated coefficients can be estimated via the bootstrap. We compare UniReg to least squares in the diabetes example of Section 4, and study it in more detail in Section 8.1.
@@ 2.3 Two examples
For the one realization from the same setup but taking $n=300,p=100$, Figure 2 shows the number of nonzero coefficients for lasso and uniLasso as a function of $\lambda$. The right end of the plot represents least squares (for the lasso) and uniReg (for uniLasso). Least squares of course uses all $p=100$ features for $\lambda=0$, but remarkably uniReg uses fewer than 20. It seems that the LOO feature scaling in step 1 and non-negative constraints in step 2 provide implicit regularization.
Figure 2: Lasso and uniLasso: size of support set as $\lambda\rightarrow 0$ for low-SNR setting with $n=300,p=100$.
@@ 6 Theoretical analysis
Let $\delta_{j,k}$ denote the population value of the univariate regression of $X_{j}$ on $X_{k}$. Suppose that for every pair of distinct indices $j,k\in S$, $\delta_{j,k}\geq 0$ if $\gamma_{j}$ and $\gamma_{k}$ have the same sign, and $\delta_{j,k}\leq 0$ if $\gamma_{j}$ and $\gamma_{k}$ have opposite signs. Then $\beta_{j}$ is nonzero and has the same sign as $\gamma_{j}$ for each $j\in S$. Moreover, $|\beta_{j}|\geq|\gamma_{j}|$ for each $j\in S$.
The last result, partly due to Ryan Tibshirani, gives a necessary and sufficient condition for equality of signs.
Theorem 3.
Suppose that $Y=\sum_{j\in S}\gamma_{j}X_{j}+\epsilon$, where all the $\gamma_{j}$’s are nonzero, and $\epsilon$ has zero mean, finite variance, and is independent of $(X_{j})_{j\in S}$. Let $\beta_{j}$ be the coefficient of $X_{j}$ in the univariate (population) regression of $Y$ on $X_{j}$. Let $\delta_{kj}$ be the coefficient of $X_{j}$ in the univariate (population) regression of $X_{k}$ on $X_{j}$. For each $j$, let $A_{j}$ be the set of $k\in S$ for which $\operatorname{sign}(\delta_{kj})=\operatorname{sign}(\gamma_{k}\gamma_{j})$ or $\delta_{kj}=0$. Then $\beta_{j}\gamma_{j}\geq 0$ if and only if
$\sum_{k\notin A_{j}}|\gamma_{k}\delta_{kj}|\leq\sum_{k\in A_{j}}|\gamma_{k}\delta_{kj}|.$
In particular, if $A_{j}=S$, then this holds.
@@ 7 Some simulation results
8 The unregularized setting: uniReg
8.1 Simulated data
Here we consider the case where $\lambda=0$, so that there is no $\ell_{1}$ penalty, as detailed in equation (4). While the main use case has $n>p$, we note that uniReg can be defined for $p>n$ by taking $\lambda\rightarrow 0$ in uniLasso. Computationally, this will approximately give the solution with minimum $\ell_{1}$ norm, in the same way that glmnet gives the least squares solution with minimum $\ell_{1}$ norm.
Conveniently even without the $\ell_{1}$ penalty, the non-negativity constraint still promotes sparsity in the solution. In Table 6 we compare least squares with uniReg in a subset of the simulated settings defined in Section 7. UniReg performs remarkably well, often winning in MSE and yielding consistently sparser models.
Table 6: Comparison of least squares with uniReg in a subset of simulated settings defined earlier.
$N=300,p=100$
MSE
setting family n p SNR mse-LS mse-uniReg
1 low-SNR gaussian 300 100 0.318 97.678 69.655
2 medium-SNR gaussian 300 100 1.272 24.428 18.559
3 high-SNR gaussian 300 100 5.087 6.104 5.82
4 homecourt gaussian 300 100 0.228 149.683 104.979
Support
setting family n p SNR supp-LS supp-uniReg
1 low-SNR gaussian 300 100 0.318 100 8
2 medium-SNR gaussian 300 100 1.272 100 9
3 high-SNR gaussian 300 100 5.087 100 7
4 homecourt gaussian 300 100 0.228 100 13
$N=300,p=30$
MSE
setting family n p SNR mse-LS mse-uniReg
1 low-SNR gaussian 300 30 0.092 71.339 65.171
2 medium-SNR gaussian 300 30 0.369 17.836 16.489
3 high-SNR gaussian 300 30 1.475 4.459 4.178
4 homecourt gaussian 300 30 0.059 111.771 101.542
Support
setting family n p SNR supp-LS supp-uniReg
1 low-SNR gaussian 300 30 0.092 30 4
2 medium-SNR gaussian 300 30 0.369 30 4
3 high-SNR gaussian 300 30 1.475 30 4
4 homecourt gaussian 300 30 0.059 30 4
$N=300,p=1000$
MSE
setting family n p SNR mse-LS mse-uniReg
1 low-SNR gaussian 300 1000 0.578 842.191 510.837
2 medium-SNR gaussian 300 1000 1.48 374.915 249.076
3 high-SNR gaussian 300 1000 3.613 186.171 147.276
4 homecourt gaussian 300 1000 1.53 577.852 420.508
Support
setting family n p SNR supp-LS supp-uniReg
1 low-SNR gaussian 300 1000 0.578 1000 24.5
2 medium-SNR gaussian 300 1000 1.48 1000 27
3 high-SNR gaussian 300 1000 3.613 1000 27
4 homecourt gaussian 300 1000 1.53 1000 135.5
8.2 Real data
We compared least squares and uniReg on the regression datasets from the UCI database. There were 23 datasets in all. Some had multiple targets: we treated each target separately and averaged the results. We randomly sampled each dataset into a $70/30$ training/test split, and report the test set MSE and support in Figure 10. One dataset had s small sample ($n=60$) which produced a very large MSE from least squares; we omit this from the plot for visibility.
Figure 10: LS vs uniReg on regression datasets from the UCI database. All datasets have $n>p$. The left plot shows the MSE ratio on the log scale. The right panel shows the support sizes of the chosen models.
The two methods perform similarly in MSE; but uniLasso has much smaller support, averaging about $1/3$ that of the least squares support (which has the number of features $p$).
9 Real data examples for uniLasso
@@ 10 More examples
10.1 UniLasso and adversarial attacks
In machine learning terminology, an adversarial attack creates an altered input that leads to a large error in the model’s prediction. In image recognition problems, the attack also produces an image that is imperceptible to humans, In statistical terms, an adversarial attack creates features in the test data that come from a different distribution than the training data,
A common way of creating an adversarial attack is to add noise in the direction of the gradient of the objective function (e.g. ?). We investigate this here in the context of the proteomics example of Table 1. We carried out the following process:
(a) Randomly split the data into a 70/30 train test split.
(b) Fit lasso and uniLasso to the training data
(c) Add varying levels of Gaussian noise to the test features in the direction of the (sub) gradient of each model from (b), and compute the test set AUC.
We repeated this 200 times and plot the results in Figure 12.
Figure 12: Results of adversarial attacks in the proteomics example of Table 1: test set AUC
We see that uniLasso is less affected by the attack than the lasso, presumably because the univariate guidance disallows sign changes and the individual features that are selected are stronger on their own.
@@ 10.2 Multiclass setting and small round blue cell cancers
mean errors out of 25 sd mean(supp) sd(supp)
@@ 15 Discussion
Draft versions of the UniLasso package can be found at
The Python package was written by Sophia Lu.
@@ Acknowledgments
The authors would like to thank Jerome Friedman, Alden Green, Sam Gross, Lucas Janson, Mert Pilanci, Jonathan Taylor, Ryan Tibshirani and Hui Zou for helpful discussions. We especially thank Ryan for posing and proving theorem 3., and Alden for the suggestion to investigate adversarial attacks in Section 10. S.C. was supported by NSF grants DMS-2113242 and DMS-2153654. T.H was supported by grant 5R01 EB 001988-21 from the National Institutes of Health; R.T. was supported by the NIH (5R01EB001988- 16) and the NSF (19DMS1208164).
@@
Proof of Theorem 3
Proof.
Note that
$\beta_{j}$ $=\frac{\mathrm{Cov}(Y,X_{j})}{\mathrm{Var}(X_{j})}$
$=\sum_{k\in S}\gamma_{k}\frac{\mathrm{Cov}(X_{k},X_{j})}{\mathrm{Var}(X_{j})}$
$=\sum_{k\in S}\gamma_{k}\delta_{kj}.$
Thus,
$\frac{\beta_{j}}{\gamma_{j}}$ $=\sum_{k\in S}\frac{\gamma_{k}}{\gamma_{j}}\delta_{kj}$
$=\sum_{k\in A_{j}}\biggl{|}\frac{\gamma_{k}}{\gamma_{j}}\delta_{kj}\biggr{|}-\sum_{k\notin A_{j}}\biggl{|}\frac{\gamma_{k}}{\gamma_{j}}\delta_{kj}\biggr{|}$
$=\frac{\sum_{k\in A_{j}}|\gamma_{k}\delta_{kj}|-\sum_{k\notin A_{j}}|\gamma_{k}\delta_{kj}|}{|\gamma_{j}|}.$
Since $\beta_{j}\gamma_{j}\geq 0$ if and only if $\beta_{j}/\gamma_{j}\geq 0$, this proves the claim. ∎